# Segre forms of singular metrics on vector bundles and Lelong numbers

### v1

@@ 2.1. Qpsh functions with analytic singularities
Let $Y$ be a complex manifold. Recall that a real valued function $u$ is qpsh (quasi-plurisubharmonic) if locally $u=v+a$, where $v$ is psh and $a$ is smooth. As already mentioned, $u$ has analytic singularities if locally
@@
where $b$ is bounded.
@@
Assume that $W$ is a connected complex manifold. If $\tau\colon W\to Y$ is holomorphic and $u$ is qpsh with (integral) analytic singularities on $Y$, then $\tau^{*}u$ is qpsh with (integral) analytic singularities in $W$. Possibly $\tau^{*}u$ is identically $\infty$- ∞.
@@
2.2. A regularization of locally bounded qpsh functions
In this subsection, we will consider a modification of the standard local regularization of psh functions through convolution, which will be suitable for approximating qas metrics with empty unbounded locus.
Proposition 2.3.
Let $\psi(x,y)$ be a locally bounded qpsh function on $U\subset{\mathbb{C}}^{n}_{x}\times{\mathbb{C}}^{m}_{y}$, assume that $\psi$ is smooth as a function of $y$, with $x$ fixed, and let $v(x,y)$ be a smooth psh function. Let $\rho_{\epsilon}(x)=\rho(x/{\epsilon})/{\epsilon}^{n}$ be a family of radial mollifiers with respect to the variable $x$. Define
where $\psi*_{x}\rho_{\epsilon}$ denotes the convolution with respect to $x$. For any $V\subset\subset U$, $\psi_{\epsilon}$ is smooth on $V$ for ${\epsilon}$ small enough, and if $A\gg 1$, then $\psi_{\epsilon}$ is increasing in ${\epsilon}$. If $\psi=u+g$, where $u$ is psh and $g$ is smooth, and $dd^{c}g\geq-B\omega$, where $\omega=dd^{c}(|x|^{2}+|y|^{2})$, then $dd^{c}\psi_{\epsilon}\geq-B\omega$.
By decomposing $dd^{c}\psi_{\epsilon}=(dd^{c}(\psi_{\epsilon}+B(|x|^{2}+|y|^{2}))-B\omega$, and applying the Bedford-Taylor theory of convergence of Monge-Ampère products for locally bounded psh functions, one obtains the following convergence.
Corollary 2.4.
Let $\psi$ and $\psi_{\epsilon}$ be as in Proposition 2.3. Then, for any $k\geq 0$,
(2.7) $(dd^{c}\psi_{\epsilon})^{k}\to(dd^{c}\psi)^{k}.$
Proof of Proposition 2.3.
We first prove that if $\psi=u$ is psh, then $\psi_{\epsilon}$ is also psh. We begin by assuming that $u$ is also smooth. In that case, by approximating the convolution by Riemann sums, $u_{\epsilon}$ may be uniformly approximated by $u_{{\epsilon},N}=\log(\sum_{k}e^{u_{k}+\log c_{k}}+e^{v+\log(A{\epsilon})})$, where $c_{k}$ are positive constants, and $u_{k}(x,y)=u(x-x_{k},y)$. Since all the functions in the exponents are psh, $u_{{\epsilon},N}$ is also psh, and thus so is $u_{\epsilon}$. If $u$ is psh, we may choose a sequence of smooth functions $u_{k}$ decreasing to $u$, and then $(u_{k})_{\epsilon}$ decreases to $u_{\epsilon}$, and hence $u_{\epsilon}$ is psh.
We consider now the case when $\psi=u+g$, where $g$ is smooth, and show that $dd^{c}\psi_{\epsilon}\geq-B\omega$, or equivalently, that $\psi_{\epsilon}^{\prime}=\psi_{\epsilon}+B(|x|^{2}+|y|^{2})$ is psh. As above, it suffices to prove the statement assuming that $u$ is smooth. We may write
$\psi_{\epsilon}^{\prime}(x,y)=\log((e^{\psi}*_{x}\rho_{\epsilon})e^{B(|x|^{2}+|y|^{2})}+e^{v+\log(A{\epsilon})+B(|x|^{2}+|y|^{2}}),$
and as above, $\psi_{\epsilon}^{\prime}$ may be uniformly approximated by
$\psi_{{\epsilon},N}^{\prime}=\log(\sum e^{\psi_{k}+\log c_{k}+B(|x|^{2}+|y|^{2})}+e^{v+\log(A{\epsilon})}).$
Since $\psi_{k}$ is just a translation of $\psi$, and $dd^{c}(|x|^{2}+|y|^{2})$ is translation invariant, it follows that all the functions in the exponents are psh, and thus $\psi_{{\epsilon},N}^{\prime}$ is psh, and hence so is $\psi_{\epsilon}^{\prime}$.
Note that $\psi_{\epsilon}$ is smooth as a function of $x$ and $y$. It remains to prove that it is increasing in ${\epsilon}$ on $V$, for $A\gg 1$. To see this, we may fix $y$ and consider everything as a function of $x$. We write $\psi=u+g$, where $u$ is psh and $g$ is smooth on $X$. We then decompose $e^{\psi_{\epsilon}}$ as
[…]
Since $e^{u}$ is psh, the first term of the right hand side is increasing in $\epsilon$. It then suffices to prove that the middle term is differentiable in $\epsilon$, with a derivative less than $Ae^{v}$ to see that the whole expression is increasing in $\epsilon$. To see this, note that $(d/d\epsilon)\rho_{\epsilon}(x^{\prime})={\mathcal{O}}(1/\epsilon^{n+1})$, and has support where $|x^{\prime}|\leq\epsilon$. Since $|e^{g(x-x^{\prime})}-e^{g(x)}|\leq C\epsilon$ for $|x^{\prime}|\leq\epsilon$ (with $C$ depending on the $C^{1}$-norm of $e^{g}$), we obtain that the derivative of the middle term is bounded by a constant, which is smaller than $Ae^{v}$ on $V$ if $A$ is chosen large enough. ∎
@@ 2.2. A regularization of locally bounded qpsh functions
Let $E\to X$ be a holomorphic vector bundle. As already mention in the introduction we have the following definition:
@@
We say that a real function is qas if it is qpsh with integral analytic singularities. A singular metric $h$ on $E\to X$ is qas if the function $\alpha\mapsto\log|\alpha|^{2}_{h}$ is qas on the total space $E$. In that case $(E,h)$ is said to be qas.
We will often drop the metric from the notation, and simply say that $E$ is a qas bundle. We will then also often denote the norm simply by $|\alpha|$.
@@
If $L$ is a line bundle with metric $e^{\psi}$, then $L$ is qas precisely if the representation of $\psi$ in any local frame is qas. ∎
@@
If we decompose the curvature $\Theta_{h}$ of $h$ as $\sum\Theta_{jk}dx_{j}{\wedge}d\bar{x}_{k}$, then $h$ is Griffiths negative if $\sum(\Theta_{jk}s,s)v_{j}\bar{v}_{k}\leq 0$ for any section $s$ of $E$ and any tangent vector $v=\sum v_{j}\frac{\partial}{\partial x_{j}}$. Notice that $\Theta_{h^{\prime}}=\Theta_{h}-A\omega_{x}$, where $\omega_{x}$ is the standard Kähler form on $U\subset{\mathbb{C}}^{N}_{x}$. By possibly shrinking $U$, we may choose $A$ such that $|\sum(\Theta_{jk}s,s)v_{j}\bar{v}_{k}|\leq A|s|^{2}|v|^{2}$, which implies that $h^{\prime}$ is Griffiths negative. ∎
The statement means that locally there is a negative line bundle $L_{A}\to U$ such that $E\otimes L_{A}$ is Griffiths negative. It is well-known that if $(E,h)$ is Griffiths negative, then $\log|\alpha|^{2}_{h}$ is psh on $E$, see, e.g., [9]. Thus we have
@@
(iii) If $E_{j}\to X$, $j=1,2$, are qas, then $E_{1}\oplus E_{2}\to X$ is qas.
(iv) If $E\to X$ has two qas metrics $h_{1}$ and $h_{2}$, then $h_{1}+h_{2}$ is qas. That is, $|\alpha|^{2}=|\alpha|^{2}_{h_{1}}+|\alpha|^{2}_{h_{2}}$ is qas.
@@
If $\tau\colon W\to X$, then $|\tau^{*}\alpha|^{2}_{\tau^{*}h}=\tau^{*}|\alpha|^{2}_{h}$ so (i) follows from Lemma 2.1. To see (ii), consider the holomorphic mapping $E\to F$, $(x,\alpha)\mapsto(x,g(x)\alpha)$. Since the pullback under a holomorphic mapping preserves the property qas, (ii) follows.
Locally $E_{j}=X\times{\mathbb{C}}_{\alpha_{j}}^{r_{j}}$. If $\log|\alpha_{j}|^{2}$ is qas on $E_{j}$, for $j=1,2$, then so is its pullback to $E_{1}\oplus E_{2}$. Thus $\log|\alpha_{j}|^{2}$ is qas on $E_{1}\oplus E_{2}$, so that by Lemma 2.2, $\log(|\alpha_{1}|^{2}+|\alpha_{2}|^{2})$ is qas on $E_{1}\oplus E_{2}$. We get (iv) by applying (i) and (iii) to $(E,h_{1})$, $(E,h_{2})$ and the mapping $E\to E\oplus E;\ \ \alpha\mapsto(\alpha,\alpha).$ For (v), note that if $e_{L}$ is a local frame for $L$, and $e^{\psi_{L}}$ the corresponding representation of the metric, then $\log|\alpha\otimes e_{L}|^{2}=\log|\alpha|^{2}+\psi_{L}$, so (v) follows by Lemma 2.2. ∎
@@
We say that a function is ${\mathbb{Q}}$-qas or ${\mathbb{R}}$-qas if it is qpsh with analytic singularities, and the local constants $c$ in (2.5) are in ${\mathbb{Q}}_{\geq 0}$ or ${\mathbb{R}}_{\geq 0}$, and that a vector bundle is ${\mathbb{Q}}$-qas or ${\mathbb{R}}$-qas if $\alpha\mapsto\log|\alpha|^{2}$ is of the corresponding class.
Note that Lemma 2.1 holds for ${\mathbb{Q}}$-qas and ${\mathbb{R}}$-qas functions, and consequently Lemma 3.5 (i)-(ii) holds for ${\mathbb{Q}}$-qas and ${\mathbb{R}}$-qas metrics. We also have that Lemma 2.2 holds for ${\mathbb{Q}}$-qas functions, but not necessarily for ${\mathbb{R}}$-qas functions (since $c_{1}\log|f_{1}|^{2}+c_{2}\log|f_{2}|^{2}$ and $\log(|f_{1}|^{2c_{1}}+|f_{2}|^{2c_{2}})$ do not necessarily have analytic singularities if $c_{1}$ and $c_{2}$ are linearly independent over ${\mathbb{Q}}$). Consequently, the corresponding version of Lemma 3.5 (iii)-(v) holds for ${\mathbb{Q}}$-qas metrics, but not necessarily for ${\mathbb{R}}$-qas metrics. ∎
@@
For any $x_{0}\in X$, there is a neighborhood $U\subset X$ and holomorphic $F_{k}(x,\alpha)=f_{k}(x)\cdot\alpha$ in $E|_{U}\simeq U\times{\mathbb{C}}^{r}_{\alpha}$, $k=1,\ldots,m$, such that
(3.1) $\sum_{k=1}^{m}|F_{k}(x,\alpha)|\sim|\alpha|_{h(x)}.$
Proof.
Assuming say that $x_{0}=0\in{\mathbb{C}}^{N}$ in a local embedding, by the assumption of analytic singularities, there are holomorphic functions $F_{k}(x,\alpha)$ in a set $|x|\leq c,|\alpha|\leq c,\ c>0$ such that (3.1) holds. Since $F_{k}(x,0)=0$ we have that $F_{k}(x,\alpha)=f_{k}(x)\cdot\alpha+F^{\prime}_{k}$, where $F^{\prime}_{k}$ is ${\mathcal{O}}(|\alpha|_{0}^{2})$ (where $|\alpha|_{0}^{2}$ denotes the trivial metric on ${\mathbb{C}}^{r}_{\alpha}$). We write the tuple as $F=f(x)\cdot\alpha+F^{\prime}$. For $x\leq c$, $|\alpha|=c$ and $|\lambda|\leq 1$, we have
$F(x,\lambda\alpha)=\lambda f(x)\cdot\alpha+{\mathcal{O}}(|\lambda|^{2}).$
Thus
$|\lambda||\alpha|_{h(x)}\leq C|F(x,\lambda\alpha)|\leq C|\lambda||f(x)\cdot\alpha|+C|\lambda|^{2},$
so that
By homogeneity this holds for all $\alpha$. In the same way we get the opposite estimate. Thus (3.1) holds with $f(x)\cdot\alpha$ instead of $F(x,\alpha)$. However, $f(x)\cdot\alpha$ is defined in the whole set $U\times{\mathbb{C}}^{r}$ and (3.1) holds there by homogeneity. ∎
Thus at least locally in $x$, $|\cdot|_{h}^{2}$ is comparable with a metric as in Example 1.1.
Proposition 3.8.
If $(E,h)$ is qas, then $(L_{E},e^{\psi})$ is qas.
Proof.
Indeed, by cf. (2.2), $\psi$ is qas on ${\mathbb{P}}(E)$ if $\log|\alpha|_{h}^{2}$ is qas on $E$, or in fact even on $E\setminus 0_{E}$. ∎
Notice by the way that $E$ is Griffiths negative, i.e., $\psi$ is psh, if and only if $e^{\psi}$ is a negatively curved metric.
Remark 3.9.
Almost all results in this paper only depend on that $L_{E}$ is qas. At a few places, Lemma 3.5 (ii)-(iii) and Lemma 3.7, and the part of Theorem 1.4 that the multiplicities are non-negative, we need the (possibly) stronger assumption that $E$ is qas. The remainder of the results in the article which are formulated for qas vector bundles $E$ indeed hold if $L_{E}$ is assumed to be qas.
In [18], the standing assumption is that $E$ is equipped with a metric $h$ such that the induced metric on $L_{E}$ is psh with analytic singularities. Note that by [18, Proposition 5.4], this is equivalent to $\log|\alpha|_{h}^{2}$ being psh with analytic singularities on $E\setminus 0_{E}$. The same arguments yield the corresponding statement when psh is replaced by qpsh, i.e., for qas metrics and functions. For such metrics, similarly to [18, Proposition 5.2], one may show that if the metric on $L_{E}$ is qas, then $\log|\alpha|_{h}^{2}$ is in fact qpsh on all of $E$. However we do not know if $\log|\alpha|_{h}^{2}$ has analytic singularities over $0_{E}$, so it is indeed possibly a stronger assumption that $E$ is qas.
The results mentioned above where we need to assume that $E$ is qas do not appear in [18]. However, if $L_{E}$ has a psh metric with analytic singularities, then the Lelong numbers of Segre forms are non-negative. Indeed, since the Lelong numbers do not depend on the choice of reference metric, we may may locally choose a negatively curved reference metric, and then $s_{k}(E,h,h_{0})$ is a positive $(k,k)$-current, and hence it has non-negative Lelong numbers.
If the following generalization of Lemma 3.3 to the case when $L_{E}$ is qas holds, then it would imply that $E$ may be equipped with a singular metric $h^{\prime}$ comparable to $h$ which is such that induced metric on $L_{E}$ is psh with analytic singularities, and in particular, $s_{k}(E,h,h_{0})$ would have non-negative Lelong numbers also under this assumption. The generalization of Lemma 3.3 is as follows: If $E$ is equipped with a singular metric $h$ such that $L_{E}$ is qas, then $he^{A|x|^{2}}$ is a Griffiths negative singular metric in a neighborhood of $x_{0}$ for $A$ large enough, i.e., the induced metric on $L_{E}$ is psh. We do not know whether it holds. ∎
4. Generalized cycles and quasi-cycles
Let $Y$ be a complex manifold or a reduced analytic space. In [5] was introduced the ${\mathbb{Z}}$-module ${\mathcal{GZ}}(Y)$ of generalized cycles, which are currents on $Y$, and which is an extension of the module ${\mathcal{Z}}(Y)$ of analytic cycles, that is, there is an injective mapping $i\colon{\mathcal{Z}}(Y)\to{\mathcal{GZ}}(Y)$, if the elements in ${\mathcal{Z}}(Y)$ are identified with their associated Lelong currents. It was shown that many properties of ${\mathcal{Z}}(Y)$ such as unique decomposition into irreducible components, well-defined integer multiplicities at each point etc., hold for ${\mathcal{GZ}}(Y)$ as well.
Recall that if $Y$ is a reduced analytic space, then smooth forms on $Y$ are defined in terms of restrictions of smooth forms in local embeddings, cf. e.g. [2], and currents as the appropriate dual of smooth compactly supported forms.
Let ${\mathcal{QZ}}(Y)$ be the $\mathbb{Z}$-module of locally finite sums of currents of the form $\tau_{*}\gamma$, where $\tau\colon W\to Y$ is a proper holomorphic mapping, $W$ is a complex manifold, and $\gamma$ is a current on $W$ of the form
(4.1) $\gamma=dd^{c}b_{1}\wedge\cdots\wedge dd^{c}b_{p},$
where $e^{b_{j}}$ are metrics on holomorphic line bundles $L_{j}$ over $W$ such that $b_{j}$ are locally bounded and qpsh. We say that the elements in ${\mathcal{QZ}}(Y)$ are quasi-cycles on $Y$.
Remark 4.1.
If we assume all $b_{j}$ are smooth we get precisely the module ${\mathcal{GZ}}(Y)$, cf.  [5, Remark 3.1], which is thus a submodule of ${\mathcal{QZ}}(Y)$. ∎
Remark 4.2.
One can just as well consider the corresponding ${\mathbb{Q}}$-module or ${\mathbb{R}}$-module that we denote by ${\mathcal{QZ}}_{\mathbb{Q}}(Y)$ and ${\mathcal{QZ}}_{\mathbb{R}}(Y)$, respectively. It is natural to refer to the elements as rational and real quasi-cycles, respectively. ∎
The product (4.1), defined by Bedford and Taylor, is a closed current of order 00. It follows that any $\mu\in{\mathcal{QZ}}(Y)$ is a closed current of order 00. In particular, such currents may be multiplied with characteristic functions ${\bf 1}_{V}$, if $V$ is an analytic subset of $W$. If $\tau\colon Y^{\prime}\to Y$ is a proper map, and $\mu\in{\mathcal{QZ}}(Y^{\prime})$, then $\tau_{*}\mu\in{\mathcal{QZ}}(Y)$.
Most results in ${\mathcal{GZ}}(Y)$ from [5] extend to ${\mathcal{QZ}}(Y)$, with the same arguments, because of the following three properties of mixed Monge-Ampère products of the form (4.1):
Lemma 4.3.
@@ 4. Generalized cycles and quasi-cycles
Let $e^{b_{1}},\ldots,e^{b_{k}}$ be metrics on line bundles $L_{1},\dots,L_{k}$ on a complex manifold $W$ such that $b_{1},\dots,b_{k}$ are locally bounded qpsh.
@@
[…]
Here $c_{1}(L_{j})$ denotes the Chern class of $L_{j}$, which can be represented by the first Chern forms associated with any smooth Hermitian metric on $L_{j}$.
@@
It is well-known from the theory of Monge-Ampère products due to Bedford-Taylor, [8], that such products do not charge pluripolar sets, cf. also, e.g., [13, Proposition III.3.11], so in particular (4.2) holds. The equality (4.3) is a local statement and follows by a suitable regularization, see, [4, Lemma 3.4]. The last equality, (4.4), is well-known in case all $b_{j}$ are smooth. Here is a sketch of an argument in our case. If $\beta_{j}$ are smooth metrics, then $b_{j}-\beta_{j}$ are global functions. We prove by induction that (4.4) holds if $b_{1},\ldots,b_{\ell}$ are replaced by $\beta_{1},\ldots\beta_{\ell}$ for $1\leq\ell\leq k$. Assume it is true for $\ell-1$. Then by the commutativity of Monge-Ampère products it is enough to verify, with $u=b_{\ell}-\beta_{\ell}$, that

### v2

@@ 2.1. Qpsh functions with analytic singularities
Let $Y$ be a complex manifold. Recall that a real-valued function $u$ is qpsh (quasi-plurisubharmonic) if locally $u=v+a$, where $v$ is psh and $a$ is smooth. As already mentioned, $u$ has (integral) analytic singularities if locally
@@
where $b$ is bounded and $c$ is a positive (integral) number.
@@
Assume that $W$ is a connected complex manifold. If $\tau\colon W\to Y$ is holomorphic and $u$ is qpsh with (integral) analytic singularities on $Y$, then $\tau^{*}u$ is qpsh with (integral) analytic singularities in $W$. Possibly $\tau^{*}u$ is identically -$\infty$.
@@
We say that a function is qas if it is qpsh and has integral analytic singularities. We can assume then that $c=1$ in (2.5).
@@
A singular metric $h$ on a holomorphic vector bundle $E\to X$ is qas if the function $\alpha\mapsto\log|\alpha|^{2}_{h}$ is qas on the total space $E$. In that case $(E,h)$ is said to be qas. The degeneracy locus of $h$ is the set $Z\subset X$ of points $x$ such that $h(x)$ is degenerate.
Notice that a qas metric $h$ is non-degenerate if and only if $\alpha\mapsto\log|\alpha|^{2}_{h}$ is locally bounded on $E\setminus 0_{E}$.
@@
Let $L$ be a line bundle with metric $h=e^{\psi}$. We claim that $(L,e^{\psi})$ is qas if and only if $\psi$ is qas on $X$. Given a local frame, $L=U\times{\mathbb{C}}_{\alpha}$ and $\log|\alpha|^{2}_{h(x)}=\log|\alpha|^{2}+\psi(x)$. If this function is qas on $L$, then by Lemma 2.1 its pullback to $U=U\times\{1\}$, i.e., $\psi$, is qas. Conversely, if $\psi(x)$ is qas  then $\log|\alpha|^{2}+\psi(x)$ is qas on $U\times{\mathbb{C}}_{\alpha}$, cf. Lemmas 2.1 and $2.$ ∎
Sometimes we suppress $h$ and say that $E$ is a qas bundle, and write $|\alpha|$ rather than $|\alpha|_{h}$.
Lemma 3.3.
(iii) If $(E_{j},h_{j})$ are qas bundles over $X$, $j=1,2$, then $(E_{1}\oplus E_{2},h_{1}\oplus h_{2})$ is qas.
(iv) If $E\to X$ has two qas metrics $h_{1}$ and $h_{2}$, then $h_{1}+h_{2}$ is a qas metric on $E$.
Proof.
If $\tau\colon W\to X$, then $|\tau^{*}\alpha|^{2}_{\tau^{*}h}=\tau^{*}|\alpha|^{2}_{h}$ so (i) follows from Lemma 2.1. Part (ii) also follows from Lemma 2.1 by considering the holomorphic mapping $E\to F$, $(x,\alpha)\mapsto(x,g(x)\alpha)$.
Locally $E_{j}=X\times{\mathbb{C}}_{\alpha_{j}}^{r_{j}}$. If $\log|\alpha_{j}|^{2}$ is qas on $E_{j}$, $j=1,2$, then so is its pullback to $E_{1}\oplus E_{2}$. By Lemma 2.2, $\log(|\alpha_{1}|^{2}+|\alpha_{2}|^{2})$ is qas on $E_{1}\oplus E_{2}$. Thus (iii) holds. We get (iv) by applying (ii) and (iii) to $(E,h_{1})$, $(E,h_{2})$ and the mapping $E\to E\oplus E,\ \alpha\mapsto(\alpha,\alpha).$ For (v), note that if $e_{L}$ is a local frame for $L$ and $e^{\psi_{L}}$ the corresponding representation of the metric, then $\log|\alpha\otimes e_{L}|^{2}=\log|\alpha|^{2}+\psi_{L}$, so (v) follows from Lemma 2.2. ∎
Remark 3.4.
We say that a function is $\mathbb{Q}$${\mathbb{R}}-qas$ if it is qpsh with analytic singularities, and the local positive constants $c$ in (2.5) are rational/real. We say that a vector bundle is $\mathbb{Q}$${\mathbb{R}}-qas$ if $\alpha\mapsto\log|\alpha|^{2}$ is $\mathbb{Q}$${\mathbb{R}}-qas$. Lemma 2.1 holds for $\mathbb{Q}$}-qas and $\mathbb{R}$}-qas functions, and consequently Lemma 3.3 (i)-(ii) holds for $\mathbb{Q}$}-qas and $\mathbb{R}$}-qas bundles. Lemma 2.2 also holds for $\mathbb{Q}$}-qas functions, but not for $\mathbb{R}$}-qas functions in general. In fact, $\log(|f_{1}|^{2c_{1}}+|f_{2}|^{2c_{2}})$ does not have analytic singularities if $c_{1},c_{2}$ are arbitrary real numbers. Therefore, the analogues of Lemma 3.3 (iii)-(v) hold for $\mathbb{Q}$}-qas bundles, but not necessarily for $\mathbb{R}$}-qas bundles. ∎
@@
If the curvature $\Theta_{h}$ of $h$ is $\sum\Theta_{jk}dx_{j}{\wedge}d\bar{x}_{k}$, then $h$ is Griffiths negative if $\sum(\Theta_{jk}s,s)v_{j}\bar{v}_{k}\leq 0$ for any section $s$ of $E$ and tangent vector $v=\sum v_{j}\frac{\partial}{\partial x_{j}}$. Notice that $\Theta_{h^{\prime}}=\Theta_{h}-I_{E}A\omega_{x}$, where $\omega_{x}$ is the standard Kähler form on $U\subset{\mathbb{C}}^{N}_{x}$ and $I_{E}$ is the identity endomorphism on $E$. By possibly shrinking $U$, we may choose $A$ such that $|\sum(\Theta_{jk}s,s)v_{j}\bar{v}_{k}|\leq A|s|^{2}|v|^{2}$. Then $h^{\prime}$ is Griffiths negative. ∎
The statement means that locally there is a negative trivial line bundle $L_{A}\to U$ such that $E\otimes L_{A}$ is Griffiths negative. It is well-known that if $(E,h)$ is Griffiths negative, then $\log|\alpha|^{2}_{h}$ is psh on $E$, see, e.g., [9]. Thus we have
@@
Assume that $h$ is a qas metric. For any $x_{0}\in X$ there is a neighborhood $U\subset X$, holomorphic $r$-tuples $f_{k}(x)$, $k=1,\ldots,m$, and $C>0$, such that in $E|_{U}\simeq U\times{\mathbb{C}}^{r}_{\alpha}$,
(3.1) $(1/C)|\alpha|_{h(x)}\leq\sum_{k=1}^{m}|f_{k}(x)\cdot\alpha)|\leq C|\alpha|_{h(x)}.$
@@
Let $x_{0}=0\in{\mathbb{C}}^{n}$ and let $|\alpha|$ be the standard norm on ${\mathbb{C}}^{r}_{\alpha}$. The assumption that $\log|\alpha|^{2}_{h(x)}$ has analytic singularities means that there are holomorphic functions $F_{k}(x,\alpha)$ in a set $|x|\leq c,|\alpha|\leq c,\ c>0$, such that (3.1) holds there with $F_{k}(x,\alpha)$ instead of $f_{k}(x)\cdot\alpha$. Since $F_{k}(x,0)=0$ we have that $F_{k}(x,\alpha)=f_{k}(x)\cdot\alpha+F^{\prime}_{k}$, where $F^{\prime}_{k}$ is ${\mathcal{O}}(|\alpha|^{2})$. We write the tuple as $F=f(x)\cdot\alpha+F^{\prime}$. For $x\leq c$, $|\alpha|=c$ and $|\lambda|\leq 1$, $F(x,\lambda\alpha)=\lambda f(x)\cdot\alpha+{\mathcal{O}}(|\lambda|^{2}).$ Thus
$|\lambda||\alpha|_{h(x)}\leq C|F(x,\lambda\alpha)|\leq C|\lambda||f(x)\cdot\alpha|+C|\lambda|^{2}.$
Letting $\lambda\to 0$ we get
By homogeneity this holds for all $\alpha$. In the same way we get the opposite inequality. Thus (3.1) holds with $f(x)\cdot\alpha$ instead of $F(x,\alpha)$. Moreover, $f(x)\cdot\alpha$ is defined in the whole set $U\times{\mathbb{C}}^{r}$ and (3.1) holds there by homogeneity. ∎
In $U$ thus $|\cdot|_{h}^{2}$ is comparable to the metric in Example 1.1, if $F=U\times{\mathbb{C}}^{m}$ with trivial metric, and $g=(f_{1},\ldots,f_{m})$.
Proposition 3.8.
If $E\to X$ is qas, then $L_{E}\to{\mathbb{P}}(E)$ is qas.
Proof.
By Lemma 3.3 (i), $p^{*}E\to{\mathbb{P}}(E)$ is qas, and by (ii) its subbundle $L_{E}$ is qas. ∎
@@
Almost all results in this paper assuming that $E$ is qas only depend on that $L_{E}$ is qas. At a few places, Lemma 3.3 (ii)-(iii) and Lemma 3.7, and the non-negativity of multiplicities in Theorem 1.3, we use the (possibly) stronger assumption that $E$ is qas. Also if $L_{E}$ has a psh metric with analytic singularities, then the multiplicities are non-negative. Indeed, locally we can choose a negative reference metric; then $s_{k}(E,h,h_{0})$ is a positive $(k,k)$-current, and hence it has non-negative Lelong numbers. In [18], the standing assumption is that $E$ is equipped with a metric $h$ such that $L_{E}$ is psh with analytic singularities. By the arguments in the proof of [18, Proposition 5.4] it follows that $L_{E}$ is qas if and only if $\log|\alpha|_{h}^{2}$ is qas on $E\setminus 0_{E}$. Following the proof of [18, Proposition 5.2], one can show that if $L_{E}$ is qas, then $\log|\alpha|_{h}^{2}$ is qpsh on all of $E$. However we do not know if $\log|\alpha|_{h}^{2}$ has analytic singularities over $0_{E}$, so $E$ being qas is possibly a stronger assumption, unless ${\rm rank\,}E=1$, cf. Example 3.2. ∎
4. Generalized cycles and quasi-cycles
Let $Y$ be a complex manifold or a reduced analytic space. In [5] was introduced the $\mathbb{Z}$}-module ${\mathcal{GZ}}(Y)$ of generalized cycles, which are closed currents on $Y$. This module is an extension of the module ${\mathcal{Z}}(Y)$ of analytic cycles, i.e., there is an injective mapping $i\colon{\mathcal{Z}}(Y)\to{\mathcal{GZ}}(Y)$, if the elements in ${\mathcal{Z}}(Y)$ are identified with their associated Lelong currents. It was shown that many properties of ${\mathcal{Z}}(Y)$ such as unique decomposition into irreducible components, well-defined integer multiplicities at each point etc., hold for ${\mathcal{GZ}}(Y)$ as well.
Recall that if $Y$ is a reduced analytic space, then smooth forms on $Y$ are defined in terms of restrictions of smooth forms in local embeddings, see, e.g., [2], and currents are dual elements as in the smooth case.
Let ${\mathcal{QZ}}(Y)$ be the $\mathbb{Z}$-module of locally finite sums of currents of the form $\tau_{*}\gamma$, where $\tau\colon W\to Y$ is a proper holomorphic mapping, $W$ is a complex manifold, and $\gamma$ is a Bedford-Taylor product, [8],
(4.1) $\gamma=dd^{c}b_{1}\wedge\cdots\wedge dd^{c}b_{p},$
on $W$, where $e^{b_{j}}$ are metrics on holomorphic line bundles $L_{j}$ over $W$ such that $b_{j}$ are locally bounded and qpsh. We say that the elements in ${\mathcal{QZ}}(Y)$ are quasi-cycles on $Y$.
Remark 4.1.
If we assume all $b_{j}$ be smooth we get precisely the module ${\mathcal{GZ}}(Y)$, cf. [5, Remark 3.1], which is thus a submodule of ${\mathcal{QZ}}(Y)$. ∎
Remark 4.2.
One can just as well consider the corresponding $\mathbb{Q}$}-module or $\mathbb{R}$}-module that we denote by ${\mathcal{QZ}}_{\mathbb{Q}}(Y)$ and ${\mathcal{QZ}}_{\mathbb{R}}(Y)$, respectively. ∎
The product (4.1) is a closed current of order 0 so it follows that any $\mu\in{\mathcal{QZ}}(Y)$ is a closed current of order 0. If $\tau\colon Y^{\prime}\to Y$ is a proper map and $\mu\in{\mathcal{QZ}}(Y^{\prime})$, then $\tau_{*}\mu\in{\mathcal{QZ}}(Y)$.
Most results about ${\mathcal{GZ}}(Y)$ in [5] extend to ${\mathcal{QZ}}(Y)$, with the same arguments, because of the following three properties of mixed Monge-Ampère products of the form (4.1):
@@ 4. Generalized cycles and quasi-cycles
Let $e^{b_{1}},\ldots,e^{b_{k}}$ be metrics on line bundles $L_{1},\dots,L_{k}$ on a complex manifold $W$ such that $b_{1},\dots,b_{k}$ are locally bounded and qpsh.
@@
(4.4) $\int_{W}dd^{c}b_{1}\wedge\cdots\wedge dd^{c}b_{k}=\int_{W}c_{1}(L_{1})\wedge\cdots\wedge c_{1}(L_{k}).$
Here $c_{1}(L_{j})$ denotes the Chern class of $L_{j}$, which can be represented by the first Chern form associated with any smooth Hermitian metric on $L_{j}$.
@@
It is well-known that products like (4.1), do not charge pluripolar sets, see, e.g., [13, Proposition III.3.11]. In particular (4.2) holds. The equality (4.3) is a local statement and follows by a suitable regularization, see, e.g., [4, Lemma 3.4]. The last equality, (4.4), is well-known in case all $b_{j}$ are smooth. Here is a sketch of an argument in our case. If $\beta_{j}$ are smooth metrics, then $b_{j}-\beta_{j}$ are global functions. We prove by induction that (4.4) holds if $b_{1},\ldots,b_{\ell}$ are replaced by $\beta_{1},\ldots\beta_{\ell}$ for $1\leq\ell\leq k$. Assume it is true for $\ell-1$. Then by the commutativity of Monge-Ampère products it is enough to verify, with $u=b_{\ell}-\beta_{\ell}$, that